A step by step example on how to go about this question would be great!

Given a Turing machine M, consider the following language LLEFT {(( Describe a Turing machine that decides this language. M), r)| M moves its head to the left at least once during its execution on input x}.
Answer #1

ANS:

IF YOU LIKE THE ANSWER PLZ GIVE THUMBSUP IT WILL HELP ME ALOT

A turing machine, which moves its head at least once during its execution on input x, can be said as, machine takes input x, moves its head to left and can move futher left without even immediate x.

So we describe Turing machine:

Input alphabet \sum =\left \{ x \right \}

tape alphabet \Gamma =\left \{ x,\Lambda \right \} where \Lambda is blank symbol

q_{s} is start state

The transition function \delta is specified by following instruction:

1) q_{s}\Lambda \rightarrow q_{s}\Lambda R

2) q_{s}x \rightarrow q_{0}x N

3) q_{0} x\rightarrow q_{1}xL

4) q_{1} x\rightarrow q_{0}xL

5) q_{2} x\rightarrow q_{0}xL

6) q_{1} \Lambda \rightarrow q_{2}\Lambda L

7) q_{2} \Lambda \rightarrow q_{1}\Lambda L

Now lets take example, say input string is \Lambda \Lambda \Lambda \Lambda x\Lambda

a) on accpeting \Lambda in initial state it will move forward till x is found and enters state q_{0}} ( by transition function1 and 2)

b) in this state, on accepting x, it will move left and enters state q_{1}} ( by transition function 3)

c) here, on accepting \Lambda, it will continue to move left and enters q_{2} ( by transition function 6)

d) here, on accepting \Lambda, it will again go to state q_{1} and so on. ( by transition function 7)

Thus the head will move atleast once when input x occurs.

If there were other input x during step b or c, it will go to state q_{0} ( by transition function 4 and 5) and will ensure again atleast one left move.